OpenStudy (anonymous):

determine whether the following pairs of lines are parallel, perpendicular, or neither 5x-6y=19 6x+5y=-30

OpenStudy (anonymous):

5x-6y=19 -6y = 19-5x y = (5/6)x-(19/6) 6x+5y=-30 5y= -30-6x y= (-6/5)x -6 slopes are 5/6 and -6/5 they are negative reciprocals, therefore the lines are perpendicular